OpenStudy (anonymous):

Solve the compound inequality: 4> 2x + 3 or 12 <= - 2x + 5

OpenStudy (anonymous):

So what does oo mean?

OpenStudy (saifoo.khan):

that is the INFINITY!

OpenStudy (anonymous):

Okay...thanks

jimthompson5910 (jim_thompson5910):

4> 2x + 3 or 12 <= - 2x + 5 4-3> 2x or 12-5 <= - 2x 1> 2x or 7 <= - 2x 1/2> x or -7/2 >= x x < 1/2 or x <= -7/2 Combine the two to get x < 1/2 So the answer is x < 1/2

OpenStudy (anonymous):

And what does the upside down U mean and the right side U mean?

OpenStudy (saifoo.khan):

U = Union.

OpenStudy (nilankshi):

rite

OpenStudy (saifoo.khan):

Let's go on that Room.